### v1

# FuSeFL: Fully Secure and Scalable Cross-Silo Federated Learning

Abstract
Federated Learning (FL) enables collaborative model training without centralizing client data, making it attractive for privacy-sensitive domains. While existing approaches employ cryptographic techniques such as homomorphic encryption, differential privacy, or secure multiparty computation to mitigate inference attacks—including model inversion, membership inference, and gradient leakage—they often suffer from high computational, communication, or memory overheads. Moreover, many methods overlook the confidentiality of the global model itself, which may be proprietary and sensitive. These challenges limit the practicality of secure FL, especially in cross-silo deployments involving large datasets and strict compliance requirements.
We present FuSeFL, a fully secure and scalable FL scheme designed for cross-silo settings. FuSeFL decentralizes training across client pairs using lightweight secure multiparty computation (MPC), while confining the server’s role to secure aggregation. This design eliminates server bottlenecks, avoids data offloading, and preserves full confidentiality of data, model, and updates throughout training. FuSeFL defends against inference threats, achieves up to 95% lower communication latency and 50% lower server memory usage, and improves accuracy over prior secure FL solutions, demonstrating strong security and efficiency at scale.
1 Introduction
Federated Learning (FL) is a distributed machine learning paradigm that enables multiple clients to collaboratively train a global model without sharing their local data [1, 2, 3, 4, 5, 6, 7]. This paradigm is particularly suited for scenarios where legal, privacy, or infrastructural constraints prevent centralized data collection [8, 5]. In FL, a central server distributes a model to the clients, who perform local training and return updates, such as gradients or model weights. These updates are aggregated to refine the global model, all while the raw data remains on client devices, preserving confidentiality and reducing transmission costs.
While FL offers clear advantages in preserving data locality, it remains vulnerable to indirect privacy breaches. In particular, submitted gradients or model updates can unintentionally encode patterns from the underlying data, enabling adversaries to extract sensitive information from these values. For example, attackers may perform model inversion to reconstruct private inputs or launch membership inference attacks to determine whether a particular data point was part of the training set [9].
To mitigate such risks, a wide range of techniques have been proposed to protect the privacy of clients’ data during training. Homomorphic encryption (HE) enables secure aggregation over encrypted updates without decryption [3, 4], but incurs high computational and memory costs, often slowing down training by several orders of magnitude [10]. Differential privacy (DP) defends against inference attacks by adding noise to updates [5, 6], though stronger privacy budgets degrade model accuracy [10]. Secure multi-party computation (MPC) allows collaborative aggregation without exposing individual inputs [7, 11, 12, 13], but suffers from growing communication overhead as the number of clients increases [10].
Beyond data privacy, a critical yet often underappreciated concern in federated learning is the privacy of the model itself, particularly from the perspective of the model owner. In typical FL deployments, the global model, often developed through significant intellectual and financial investment, is shared with all participating clients, some of whom may be semi-honest or untrusted [14]. This exposes the model to risks such as extraction, reverse engineering, or unauthorized use. The concern becomes even more pressing in cross-silo FL, where clients are typically large institutions, such as hospitals, banks, or research labs, that operate under strict privacy regulations and contractual obligations. In such settings, the model is not only a valuable proprietary asset but also a potential vector for secondary leakage of sensitive information. Ensuring model confidentiality is therefore essential, not only to protect intellectual property but also to comply with legal, regulatory, and contractual boundaries.
Several recent works have attempted to address model privacy in federated learning, though each approach faces significant limitations that hinder its practicality, particularly in cross-silo deployments where efficiency, scalability, and regulatory compliance are critical.
Watermarking-based techniques have been proposed to address model ownership and intellectual property protection [15, 16]. These methods embed verifiable patterns, such as trigger inputs or parameter, level watermarks—into the model, allowing the owner to later demonstrate unauthorized use. However, watermarking does not prevent theft or misuse during training; it only enables detection after the fact. In domains like healthcare and finance, where compliance and security must be enforced upfront, such reactive mechanisms are insufficient.
WW-FL [17] adopts a fully offloaded training architecture in which secure multi-party computation (MPC) is performed across distributed clusters of servers. While this design offers strong security guarantees, it requires clients to upload their entire datasets to third-party servers, contradicting the foundational principle of cross-silo FL, where data should remain local to each silo. This architecture introduces significant communication overhead, as large volumes of sensitive data must be transmitted over the network, increasing both latency and memory usage. Moreover, it depends on geographically distributed compute infrastructure, which elevates deployment complexity and operational costs. In cross-silo environments, where clients such as hospitals or financial institutions already possess rich compute and storage capabilities, outsourcing training to external servers is not only inefficient but also unnecessary.
A similar line of work proposes offloading both the model and client data to centralized secure servers that perform training using cryptographic techniques [18] such as MPC [19, 20, 21, 22, 23, 24]. While these systems reduce the trust placed in individual clients, they impose substantial communication burdens due to repeated transfers of large training datasets. This issue is exacerbated in cross-silo FL, where each client holds sizable, high-value data. Additionally, centralizing sensitive datasets on untrusted servers risks violating regulatory constraints, institutional data policies, and contractual agreements, making these approaches ill-suited for regulated domains.
AriaNN [22] introduces a hybrid federated scheme that jointly protects model and data privacy by secret-sharing the model between the client and server and distributing the computation using MPC. This design prevents either party from observing the full model or training data in plaintext. While more aligned with FL principles, AriaNN suffers from serious scalability limitations. Each client must train its share of the model in coordination with the server, typically in a sequential fashion due to server-side compute constraints. Communication overhead grows linearly with client count; even modest datasets like MNIST can require tens of megabytes per client, and this cost becomes prohibitive for large datasets common in siloed environments. Furthermore, the server must maintain a dedicated model instance per client in addition to the global model, leading to unsustainable memory overhead for large-scale models such as VGG-16.
In summary, existing federated learning approaches for cross-silo settings struggle to scale beyond a few tens of clients due to excessive computational, communication, and memory overheads. Techniques like homomorphic encryption and secure model-sharing, while effective in preserving privacy, impose severe performance costs. In particular, efforts to simultaneously protect both data and model confidentiality exacerbate these challenges, rendering current solutions impractical for real-world, large-scale deployments.
To address the limitations of prior work, we propose FuSeFL, a fully secure and scalable federated learning framework tailored for cross-silo environments. FuSeFL overcomes the limitations of prior work by ensuring end-to-end confidentiality of both the data and the model, while also scaling efficiently to large client populations. The key idea is to decentralize training across client groups while confining the server’s role to lightweight secure aggregation.
In FuSeFL, clients are dynamically paired to form collaborative groups. Within each group, the global model is secret-shared between the two clients, and training is jointly executed using secure multi-party computation (MPC) over their local data. This structure prevents any party from accessing the model or data in plaintext. After training, each group submits secret-shared updates to two non-colluding aggregation servers, which perform secure aggregation without learning individual contributions.
This design resolves the critical issues in prior solutions: it avoids server-side bottlenecks, eliminates the need for data offloading, and reduces communication overhead. Moreover, by keeping both the model and updates secret-shared throughout training and aggregation, FuSeFL protects against gradient leakage, model update inference, membership inference, and model inversion attacks.
Our evaluation shows that FuSeFL (1) reduces server memory usage by nearly 50%, (2) maintains nearly constant training time as the number of clients increases, achieving linear throughput scalability, and (3) cuts communication latency by 95% compared to the state-of-the-art schemes in large-scale deployments. FuSeFL also improves model accuracy by up to 1.1% over the federated adaptation of AriaNN on MNIST.
Our key contributions are:
  • A fully secure FL scheme: We develop FuSeFL, a federated learning scheme that holistically secures model parameters, client data, and intermediate updates using lightweight MPC across client groups—addressing both model and data privacy within a unified design (§5.1).
  • Decentralized and scalable training: FuSeFL eliminates server-side training bottlenecks by distributing computation among client pairs. This design supports parallelism, avoids reliance on third-party compute infrastructure, and is well-suited for cross-silo FL environments with powerful clients (§7.2).
  • Practical efficiency with robust privacy guarantees: Our approach significantly reduces memory and communication overheads while mitigating key inference threats such as model inversion, gradient leakage, and membership inference, making FuSeFL scalable to real-world deployments involving many silos (§7.4, §7.3).
2 Motivation and Goal
2.1 Limitations of Existing Work
Despite increased attention to privacy in federated learning, fully secure FL remains constrained by fundamental system bottlenecks. We identify three key limitations:
  1. 1.
    Model Exposure and Trust Assumptions: Many existing FL schemes expose the full global model to clients during training or inference, risking model theft or misuse, particularly in regulated domains where model intellectual property is sensitive [25, 26, 15, 16]. Moreover, several privacy-preserving methods offload client data or models to trusted third-party servers using MPC or HE protocols [19, 20, 21, 22, 27, 28]. These trust assumptions violate the core principle of cross-silo FL, which requires data to remain within institutional boundaries [29, 30].
  2. 2.
    Server Bottleneck and Poor Scalability: schemes such as AriaNN-FL[22] impose significant computational load on the server. Secure training requires the server to engage individually with each client, saturating its compute resources and limiting parallelism. As shown in Fig. 1 (left axis), training time increases sharply as the number of clients grows, even when using a 16-core server. This sequential interaction model hinders scalability in larger deployments.
  3. 3.
    Communication and Storage Overheads: Secure training introduces high communication and memory demands. For example, training Network1 over MNIST with AriaNN-FL requires more than 10 MB of data exchange per client per epoch [22]. With hundreds of clients, this quickly overwhelms bandwidth limits. Additionally, the server must retain a copy of each client’s model update until aggregation is complete. For large models like VGG-16, this results in excessive memory usage: as illustrated in Fig. 1 (right axis), AriaNN-FL stores 128 full model instances when training with 128 clients.
Figure 1: Comparison of AriaNN-FL with FuSeFL, (Left vertical axis) in terms of secure training time for one epoch and (Right vertical axis) in terms of the number of model instances stored in one aggregator. All these experiments use 16 parallel computing cores for a varying number of clients. By increasing the number of clients, the training time and the number of model instances increase linearly in AriaNN-FL.
2.2 Our Goals
FuSeFL aims to deliver a practical, fully secure, and scalable cross-silo federated learning framework that safeguards both user data and the model itself, an often-overlooked dimension in secure FL literature. Designed for real-world cross-silo deployments with potentially several hundred or even thousands of clients, FuSeFL balances strict privacy guarantees with system-level scalability and efficiency. Specifically, it seeks to:
  • Ensure end-to-end confidentiality of data and models: Unlike conventional FL, where clients have full access to the global model, FuSeFL guarantees that both the model and its updates remain secret-shared and are never revealed in plaintext to any single party.
  • Remove server-side computational bottlenecks: By decentralizing local training to client groups, FuSeFL avoids the per-client training load imposed on servers in frameworks like AriaNN-FL. Servers are instead limited to lightweight secure aggregation, improving scalability and efficiency.
  • Reduce communication and memory overhead: FuSeFL forms client pairs that collaboratively train and send a single secret-shared model update per group, rather than having each client transmit updates independently. This design cuts communication volume and nearly halves server-side memory usage, improving resource efficiency in large-scale deployments.
  • Achieve secure scalability to a larger number of clients: Training is parallelized across multiple client groups, allowing FuSeFL to maintain high throughput and low latency as the number of clients grows, overcoming the scalability limitations of existing secure FL protocols.
3 Threat Model and Security Considerations
Deployment model and trust assumptions: FuSeFL targets scalable cross-silo federated learning where each client, such as a hospital or financial institution, holds a distinct set of samples over a shared feature space, a concept known as horizontal federated learning [31]. Clients are governed by regulatory frameworks and formal agreements such as HIPAA or GDPR, which limit the risk of deviation from the protocol.
The trust model assumes: (1) Clients do not collude with others in the same training group. Each may attempt to infer private information from its own view but adheres to the protocol. (2) Aggregation servers are honest-but-curious, do not collude with one another or with clients, and jointly perform MPC-based aggregation. (3) The model owner is trusted to manage protocol coordination but cannot access client data or intermediate model states.
To reduce the risk of collusion and information leakage across rounds, FuSeFL employs dynamic client grouping, reassigning clients into new training groups in each round to prevent repeated interactions. While our protocol assumes no collusion within a group during a single round, this is a well-founded assumption in cross-silo federated learning[29, 30], where clients are typically autonomous organizations with independent data silos, isolated infrastructure, and legally enforced compliance boundaries. Such structural and regulatory separation significantly reduces the feasibility of coordinated behavior required for collusion. FuSeFL is architected to support threshold secret sharing as a drop-in extension, enabling privacy preservation even with up to $t$ colluding clients per group. We plan to implement and evaluate this enhancement in future work.
Threat model: We consider a semi-honest (honest-but-curious) adversary who follows the protocol but may attempt to infer private information from observed data. The system involves three types of participants: clients, aggregation servers, and a model owner. Clients perform local training using secret-shared model parameters. Aggregation servers coordinate secure aggregation without learning the underlying values. The model owner is responsible for initializing the global model and orchestrating training (as shown in Fig. 2).
Security Guarantees: FuSeFL provides strong privacy protections through lightweight cryptographic primitives. Specifically: (1) Data confidentiality: Each client’s private data is secret-shared within its training group, ensuring that no single party can reconstruct the input. (2) Model confidentiality: The global model is secret-shared between aggregation servers and is never exposed in plaintext to any party during training or aggregation. (3) Update privacy: Local model updates are secret-shared and aggregated securely using MPC, such that individual updates remain hidden.
These guarantees mitigate risks such as unauthorized observation, gradient leakage, membership inference, model inversion, and reconstruction attacks. FuSeFL enforces these protections using efficient secure computation protocols based on additive and function secret sharing.
Out-of-scope threats: FuSeFL does not currently defend against poisoning attacks [32] and malicious participants who deviate from the protocol, submit corrupted updates, or act in a Byzantine manner. Future extensions may incorporate robust aggregation, verifiable computation, or trusted execution to address these threats. Side-channel attacks[33, 34, 35] such as timing[33] or traffic analysis[35] are also out of scope and would require separate system-level countermeasures.
Figure 2: Overview of the threat model. Each party retains private data within its trusted boundary. Adversaries may attempt to infer information about the data or model from the untrusted boundary.
4 Background
4.1 Notations and Preliminaries
Our scheme leverages MPC to preserve the privacy of unencrypted data during training. Table I summarizes the notations used throughout the paper.
TABLE I: List of Notations
Notations Explanation
$GM$ Global model in plaintext
$[GM]_{j}$ $j^{th}$ share of the secret shared global model
$LM_{i}$ $i^{th}$ local model in plaintext
$[LM_{i}]_{j}$ $j^{th}$ share of the secret shared $i^{th}$ local model
$D_{i}$ Plaintext data of the $i^{th}$ client
$[D_{i}]_{j}$ $j^{th}$ share of the secret shared data of $i^{th}$ client
4.2 Federated Learning (FL)
Traditional machine learning requires aggregating data from multiple sources onto a central server for training. However, this approach is often impractical due to privacy regulations, communication costs, and bandwidth constraints [2].
Federated Learning (FL) addresses these challenges by enabling collaborative model training across distributed clients without exposing raw data [1, 2]. As illustrated in Fig. 3, (1) the central server first distributes the current global model to a subset of participating clients. (2) Each client trains the model locally using its private dataset. (3) The clients then send their model updates back to the server. (4) The server aggregates these updates (e.g., via FedAvg [2]) to refine the global model. This process is repeated for multiple rounds until convergence.
Figure 3: Overview of Federated Learning. (1) The server sends the global model to the clients. (2) The clients then train the model locally and (3) send back their trained models to the server. (4) The server updates the global model for the next round of training.
FedAvg [2] is the most widely adopted aggregation method in federated learning. It computes the global model by taking a weighted average of the local model parameters ($w$) from participating clients:
$w_{t+1}=\sum_{k=1}^{K}\frac{n_{k}}{n}w_{t}^{k}$
Here, $w_{t}^{k}$ denotes the model parameters from client $k$ at round $t$, $n_{k}$ is the number of training samples on client $k$, and $n=\sum_{k=1}^{K}n_{k}$ is the total number of samples across all clients. This weighted averaging ensures that clients with larger datasets have proportionally greater influence on the updated global model.
Cross-silo federated learning targets deployments where each client represents a large, resource-rich organization, such as a hospital or financial institution, with stable availability and ample compute, storage, and bandwidth [29]. While these settings offer stronger infrastructure support than cross-device FL, they also present distinct challenges. The large datasets held by each silo amplify communication and memory demands, resulting in scalability bottlenecks, particularly at the aggregator, where per-client updates must be processed and stored [36, 37, 29].
4.3 Secure Multi-Party Computation (MPC)
MPC is a cryptographic technique that enables multiple parties to collaboratively compute a function over their private inputs while ensuring that the inputs remain confidential and individual shares are never disclosed during the process. By splitting sensitive data into encrypted or secret shares and processing them separately, secure computation is achieved.
MPC allows computation over encrypted data while remaining lightweight, making it a practical option for distributing computation securely across multiple clients without incurring heavy computational costs. However, as the number of participants increases, the communication overhead in MPC protocols also grows significantly.
Additive Sharing. To perform linear operations, additive sharing [38, 39, 40, 41] is a common MPC technique that splits private data into multiple secret shares, each distributed to a different party. In this approach, the data is secret-shared such that no single party can reconstruct it independently. If there are $j$ parties, a private value $X$ is represented as the sum of its shares:
$X=\sum[x]_{j}$
These shares are indistinguishable from random values and reveal no private information individually. Each party keeps its share private and does not communicate it to others. Secure computation is achieved by applying operations directly over these shares, as described in prior work [19, 39].
  • Addition: Each party performs addition locally on its share of the data. The partial results from all parties are then aggregated to compute the final result. As illustrated in Fig. 4(a), the private input $X$ is split into two shares, $X_{r}$ and $X_{c}$. After local computation, the final output $Y$ is obtained by summing the partial results $Y_{1}$ and $Y_{2}$. For $j$ parties:
    $[Z]_{j}=[X]_{j}+[Y]_{j},\quad Z=\sum\text{(locally computed)}\ [Z]_{j}$
  • Multiplication: To perform multiplication over additive shares, Beaver triples [42] are used. If $[Z]=[X]\cdot[Y]$, and the precomputed triple is $[c]=[a]\cdot[b]$ [19], each party $P_{j}$ does the following:
    • Compute local differences:
      $E_{j}=X_{j}-a_{j},\quad F_{j}=Y_{j}-b_{j}$
    • Publicly reconstruct the differences:
      $E=\sum E_{j},\quad F=\sum F_{j}$
    • Compute the local share of the final result:
      $Z_{j}=j\cdot E\cdot F+F\cdot a_{j}+E\cdot b_{j}+c_{j}$
Function Secret Sharing (FSS). Unlike additive sharing, Function Secret Sharing [43, 44, 45] splits the function itself across parties. That is, each party holds a share of the function rather than the data, and computation is done by applying these function shares to the input. If $f(X)$ is the function to evaluate and $X$ is the (possibly masked) input, then:
$f(X)=f_{1}(X)+f_{2}(X)$
Fig. 4(b) illustrates this: the private input $X$ is masked as $X_{c}=X+R$, and each party evaluates its function share on $X_{c}$. The final result is reconstructed by summing the outputs:
$f(X_{c})=f_{1}(X_{c})+f_{2}(X_{c})$
Figure 4: Examples of MPC schemes. (a) Additive sharing splits data into random shares for secure multi-party computation. (b) Function secret sharing splits the function itself across parties for privacy-preserving evaluation.
4.4 AriaNN: Federated Learning Extension
Ryffel et al. [22] extended their secure training framework to support federated learning, enabling privacy-preserving model training across distributed clients. In this FL-based variant of AriaNN, the server is treated as an untrusted third party by both clients and model owners.
The training process begins by secret-sharing the model parameters between the server and clients, each holds one share of the model. As illustrated in Fig. 5, the protocol proceeds as follows: (1) Each client secret-shares a portion of its local data with the server while retaining the other share locally. (2) Clients and the server then collaboratively train their respective shares of the local model using MPC over a number of training iterations. (3) After local training, one designated client (e.g., $C_{A}$) acts as the aggregator, collecting secret-shared updates from other clients. (4) The aggregator performs secure weighted averaging of these updates, ensuring that individual data and model privacy are preserved throughout. The server, in parallel, aggregates its corresponding shares of the model updates. (5) The updated model share $[GM]_{1}$ is redistributed privately by the aggregator to all clients for the next round, while the server also updates its copy $[GM]_{0}$ accordingly.
This process repeats for a fixed number of global epochs. At the end of training, the final model is reconstructed by combining the two secret shares:
$GM=[GM]_{0}+[GM]_{1}$
Since the model remains secret-shared throughout training, this design protects both data and model confidentiality while enabling secure collaborative learning.
Figure 5: Overview of the AriaNN-FL scheme. (2) Each client collaboratively trains its local model with the server. The server holds a share of the model and the data, while the remaining shares are securely stored on the client side. (4) $C_{A}$, i.e., client aggregator, aggregates the client-side shares of the local models.
5 Fully Secure Federated Learning (FuSeFL)
In this section, we present FuSeFL, a federated learning framework that holistically protects both data and model confidentiality while enabling scalable training in cross-silo environments. FuSeFL addresses key limitations of prior work, including privacy vulnerabilities and centralized bottlenecks. In many existing systems, the server performs per-client training or coordination, limiting parallelism and creating a performance bottleneck. This becomes especially restrictive in cross-silo settings, where clients are powerful and data-rich. FuSeFL overcomes this by decentralizing training across client groups and assigning the server only a lightweight aggregation role, enabling secure and efficient scalability.
Figure 6: Overview of FuSeFL. Clients shape groups of two. In each group, the data and model are secret shared and clients collaboratively and securely train the local model.
5.1 Overview of FuSeFL
To enable secure distributed training, FuSeFL leverages multi-party computation to keep both data and model parameters secret-shared throughout. The protocol builds on AriaNN’s MPC backend but rearchitects the system to support federated deployment, enhance scalability, and lower communication latency.
In each training round, clients are grouped into pairs. Within each group, both the model and client data ($D_{k}$ and $D_{L}$) are secret-shared to ensure that no single party can access them in plaintext. As shown in Fig. 6, the workflow proceeds as follows:
  1. 1.
    Model Distribution: The global model is secret-shared between two servers ($S_{1}$, $S_{2}$), who then transmit the shares to the two clients ($C_{k}$, $C_{l}$) in each group. Specifically, $S_{1}$ sends $[GM]_{1}$ to $C_{k}$, and $S_{2}$ sends $[GM]_{2}$ to $C_{l}$.
  2. 2.
    Secure Group Training: Clients $C_{k}$ and $C_{l}$ collaboratively train a local model over their secret-shared data ($D_{k}$ and $D_{L}$) using MPC. The model itself is maintained in a secret-shared form throughout training, ensuring confidentiality of both inputs and intermediate states.
  3. 3.
    Update Submission: After completing local training for several epochs, each client returns their share of the updated local model to the corresponding server—$C_{k}$ sends $[LM]_{1}$ to $S_{1}$ and $C_{l}$ sends $[LM]_{2}$ to $S_{2}$.
  4. 4.
    Secure Aggregation: Servers compute a weighted average (FedAvg [2]) over the secret-shared model updates. Because updates are shared across two non-colluding servers, confidentiality is preserved even during aggregation.
  5. 5.
    Iteration: The aggregated global model is redistributed to client groups for the next training round. This process repeats for multiple global epochs.
By structuring the training process around small, secure client groups and decoupling server-side training from aggregation, FuSeFL achieves a practical and scalable solution for secure FL in regulated cross-silo deployments.
5.2 Risk-Aware Dynamic Client Grouping
To reduce the likelihood of information leakage through collusion, FuSeFL employs a dynamic, adversary-aware grouping strategy in which client pairs are reshuffled every global round. This continual reconfiguration enhances robustness by minimizing repeated co-training relationships, thereby amplifying privacy through exposure diversity.
Client grouping is coordinated by the server using only public or registered metadata, such as organizational type, legal jurisdiction, or historical affiliations, while avoiding access to private data or training artifacts. This design choice supports deployment in cross-silo settings where participants (e.g., hospitals, banks, or labs) operate under distinct governance and regulatory constraints.
Each client pair is assigned a collusion risk score $\rho(C_{i},C_{j})\in[0,1]$, reflecting potential structural correlations such as shared ownership, industry overlap, or regulatory proximity. Given these scores, FuSeFL computes a set of disjoint client pairs $\mathcal{G}=\{(C_{i},C_{j})\}$ that minimize overall collusion exposure:
$\mathcal{G}^{*}=\arg\min_{\mathcal{G}\subset\binom{C}{2}}$ $\sum_{(C_{i},C_{j})\in\mathcal{G}}\rho(C_{i},C_{j})$ (1)
s.t. each $\text{each }C_{i}\text{ appears exactly once.}$.
This problem maps to a classical minimum-weight perfect matching on a complete undirected graph, where clients are nodes and edge weights correspond to collusion risk scores. The exact solution can be obtained in $O(n^{3})$ time using combinatorial optimization techniques such as the Hungarian algorithm [46]. However, for large-scale deployments, FuSeFL employs a more scalable approximation: it greedily constructs groupings by iteratively selecting the lowest-risk unpaired edge until all clients are matched. This greedy matching runs in $O(n^{2}\log n)$ and has shown strong empirical performance for similar pairwise minimization tasks [47].
Pipelined Execution: FuSeFL performs group assignment computation entirely in parallel with training. Because group matching relies solely on static metadata and does not interact with model state or local data, assignments for round $t{+}1$ are computed concurrently while training for round $t$ is ongoing. As a result, this procedure incurs no additional latency in the training workflow.
Establishing MPC Channels: Once groups are selected, each pair $(C_{i},C_{j})$ initiates a secure MPC session using authenticated metadata exchange and ephemeral key agreement. These sessions are short-lived and introduce negligible delay (typically under 30ms per pair [22]). Channels are secured via TLS or mutual-authenticated Diffie-Hellman, in alignment with enterprise-grade deployments.
This risk-aware, round-wise reshuffling reduces long-term collusion opportunities, ensures regulatory separation, and strengthens overall privacy guarantees in multi-institution FL settings.
5.3 Training Modes: Serial and Parallel
To address the trade-offs between training accuracy, latency, and memory usage, FuSeFL introduces two intra-group training modes: FuSeFL-Serial and FuSeFL-Parallel, each offering distinct performance characteristics.
In FuSeFL-Serial (Fig. 7(a)), each client group jointly trains a single model instance using the data from both clients, one after the other. First, the local model $LM_{1}$ is trained using client $C_{1}$’s data $D_{1}$ under MPC. Then, the same model is further trained using client $C_{2}$’s data $D_{2}$. This sequential training leverages the full dataset available to the group, typically leading to improved model accuracy. Since only one model instance is maintained, this mode is memory-efficient. However, it incurs higher training latency due to its serialized nature. Algorithm 1 provides implementation details.
Figure 7: (a) FuSeFL-Serial trains the same model using the data of both clients in the group. (b) FuSeFL-Parallel trains two separate local models using the data of each client.
In contrast, FuSeFL-Parallel (Fig. 7(b)) accelerates training by allowing both clients to train separate local models concurrently. The group maintains two model instances—$LM_{1}$ and $LM_{2}$—which are secret-shared among the clients. Each model is trained independently on $D_{1}$ and $D_{2}$ in parallel. After training, the clients locally aggregate the two updated models before securely transmitting the result to the servers. This mode reduces training latency and is better suited for time-sensitive deployments (evaluated in §7.2), at the cost of slightly higher memory usage and a modest trade-off in final model accuracy (evaluated in §7.1). Details are outlined in Algorithm 2).
Both training modes fully preserve data and model confidentiality through MPC, ensuring that no single client or server can reconstruct sensitive information. Table II summarizes the trade-offs between the two modes.
TABLE II: Comparison of FuSeFL-Serial and FuSeFL-Parallel in terms of performance and resource trade-offs.
Aspect FuSeFL-Serial FuSeFL-Parallel
Training Latency Higher Lower
Model Accuracy Slightly higher Slightly lower
Memory Usage Lower Higher
Best Use Case Accuracy- or memory- sensitive applications Latency-sensitive applications
Input: $j\leftarrow 2$
            Global model $[GM]_{j}$;
            Local models $[LM_{(i/2)}]_{j}$;
            Local datasets $[D_{i}]_{j}$
for s in j do
       Server s $\leftarrow$ $[GM]_{j}$
end for
for each global epoch do
      for each group g (in parallel) do
            for each client c in g do
                   for s in j do
                         Server s sends $[GM]_{j}$ to client c
                   end for
                  Client c $\leftarrow$ $[D_{c}]_{j}$
                   Client c $\leftarrow$ $[D_{(j-c)}]_{(j-c)}$
                   $[LM_{g}]_{j}$leftarrow $[GM]_{j}$
             end for
            for each local epoch do
                   clients train $[LM_{g}]$ using $[D_{0}]$
             end for
            for each local epoch do
                   clients train $[LM_{g}]$ using $[D_{1}]$
             end for
            for each client c in g do
                   for s in j do
                         client c sends $[LM_{g}]_{j}$ to server s
                   end for
             end for
       end for
      $[GM]_{j}$ = $AVG(\Sigma{[LM_{g}]_{j})})$
end for
Algorithm 1 FuSeFL-Serial
for each global epoch do
      for each group g (in parallel) do
            for each client c in g do
                   for s in j do
                         Server s sends $[GM]_{j}$ to client c
                   end for
                  Client c $\leftarrow$ $[D_{c}]_{j}$
                   Client c $\leftarrow$ $[D_{(j-c)}]_{(j-c)}$
                   $[LM_{c}]_{j}$leftarrow $[GM]_{j}$
             end for
            for each local epoch do
                   clients train $[LM_{c}]$ using $[D_{0}]$
             end for
            for each local epoch do
                   clients train $[LM_{g}]$ using $[D_{1}]$
             end for
            $[LM_{g}]=Aggregate([LM_{2n}],[LM_{2n+1}])$
       end for
      $[GM]$ = $AVG(\Sigma{[LM_{g}])})$
end for
Algorithm 2 FuSeFL-Parallel
5.4 Communication Complexity
Secure federated learning systems incur communication overhead primarily in two channels: server-to/from-client (S2C/C2S) and client-to-client (C2C).
Server-to-Client / Client-to-Server (S2C/C2S). In AriaNN-FL, the server actively participates in secure training by running MPC protocols jointly with each client across every local epoch. This introduces a severe server-side bottleneck as the number of clients grows. In contrast, FuSeFL decouples training from the server and restricts its role to model aggregation. Clients only communicate their secret-shared updates to two non-colluding servers once per global round. The overall S2C/C2S complexity is reduced to $\mathcal{O}((n/2)\cdot x)$, assuming a model size of $x$ and $n$ total clients. This significantly improves scalability and enables parallel execution. Client-to-Client (C2C). AriaNN-FL performs client-to-client communication during aggregation, where each client sends its model share to a designated aggregator for reconstruction—incurring $\mathcal{O}(n\cdot x)$ communication and requiring one trusted aggregator per round. In contrast, FuSeFL confines secure training to fixed client pairs using MPC. Each pair exchanges secret shares of data and model parameters locally, with communication parallelized across groups. This keeps per-group C2C overhead constant and independent of the total number of clients.
By restructuring the communication pattern to exploit pairwise training and two-server aggregation, FuSeFL substantially reduces both the volume and centralization of communication (evaluated in §7.3). This architectural change eliminates per-epoch server interaction, enables system-wide parallelism, and improves bandwidth efficiency without weakening security guarantees.
5.5 The Memory Complexity
Table III summarizes the number of model instances maintained by different entities under AriaNN-FL and FuSeFL (in both Serial and Parallel modes). In AriaNN-FL, the server is involved in both secure training and aggregation, requiring it to maintain $n$ local model instances—one per client—plus the global model. Similarly, the designated aggregator client stores $n$ model shares for aggregation, resulting in significant memory overhead concentrated on central parties. In contrast, FuSeFL distributes the training workload across client pairs, thereby decentralizing memory usage. Each of the two aggregation servers stores only $(n/2)+1$ model instances, cutting the server-side memory requirement by nearly 50%. On the client side, FuSeFL-Serial maintains a single model per group, yielding comparable memory usage to AriaNN-FL but with distributed storage. FuSeFL-Parallel trades memory for speed: each client trains an independent local model copy, increasing client-side memory usage to $3n$ total model instances across all clients. This redundancy enables parallel execution and accelerates training without incurring additional server load.
TABLE III: Number of instances of the model which is stored on different parties. FuSeFL almost halves the number of instances required to be stored on the server.
AriaNN-FL FuSeFL-Serial FuSeFL-Parallel
Server1 n + 1 $(n\div 2)+1$ $(n\div 2)+1$
Server2 N/A $(n\div 2)+1$ $(n\div 2)+1$
Clients n n 3n
Client Aggregator n + 1 N/A N/A
5.6 Security Guarantees
FuSeFL provides strong security guarantees for both data and model confidentiality in a distributed, partially untrusted environment. These guarantees are realized through a combination of secret sharing, secure multi-party computation (MPC), and adversary-aware system design, as defined in our threat model (Section 3). Below, we detail the key properties and the cryptographic principles underpinning them.
  • Confidentiality of Local Data, Model Parameters, and Updates: Each client retains the confidentiality of its data, model parameters, and model updates throughout the training process. FuSeFL employs additive secret sharing to ensure that no single party has access to the plaintext of any sensitive information. All values exchanged during training are masked using independently sampled random shares that are statistically indistinguishable from uniform noise. Secure evaluation of both linear and nonlinear operations is achieved using function sharing techniques as in AriaNN [22]. Since neither client nor server ever observes unmasked values, the protocol provides semantic security against honest-but-curious adversaries.
  • Correctness of Secure Training: FuSeFL builds directly on the AriaNN protocol, which guarantees the correctness of MPC-based training among two clients. Each party’s local execution follows the prescribed sequence of arithmetic and Boolean operations as encoded in the secure computation circuit. The resulting model update is functionally equivalent to that of plaintext training. For formal analysis, we refer to the proofs provided in [22].
  • Correctness of Aggregation: After each group completes training, clients submit secret shares of the locally updated model to two non-colluding aggregation servers. Each server independently performs FedAvg-style aggregation [2] over the received shares. Since weighted averaging is a linear operation, it naturally composes with additive secret sharing. Thus, the final global model obtained after reconstruction matches the federated average of all local updates. The correctness of this aggregation step is inherited from prior secure aggregation frameworks [19, 39, 40].
  • Protection Against Curious Clients: Clients never access the raw data or model parameters of their partner. Instead, they hold one share of their partner’s input and model, making reconstruction infeasible under an honest-but-curious assumption. This ensures that a semi-honest client cannot learn the counterpart’s training data or model parameters. Note that FuSeFL does not assume malicious collusion between group members; however, privacy guarantees degrade proportionally with the number of colluding parties.
5.7 Flexibility and Generalization
FuSeFL is designed to be modular with respect to the underlying secure computation protocol. While our current implementation leverages the AriaNN scheme for secure collaborative training within client groups, FuSeFL’s architecture is not bound to this specific backend. The abstraction of secure local training allows FuSeFL to seamlessly generalize to a wider class of MPC schemes.
For instance, adopting a three-party MPC protocol such as SecureNN [20] or Falcon [21] enables FuSeFL to extend its grouping strategy to triads of clients. In such a configuration, each client secret-shares its data and model with the remaining clients in the group, maintaining one share locally. Training proceeds securely over these shared inputs without revealing sensitive data to any single party. The model itself must also be secret-shared in alignment with the MPC protocol’s trust assumptions.
To maintain end-to-end confidentiality, the aggregation infrastructure must scale accordingly. Specifically, FuSeFL requires one aggregation server per client in the group to receive and process the respective secret shares. As a result, for a group of $n$ clients using an $n$-party MPC protocol, FuSeFL can tolerate up to $n{-}1$ colluding clients or servers without compromising data or model confidentiality.
This flexible design allows FuSeFL to support a broad range of MPC backends with different trust and performance trade-offs, making it suitable for deployment in diverse federated learning environments where regulatory, computational, or adversarial constraints may vary.
6 Evaluation Methodology
To assess the performance and accuracy of FuSeFL, we adopt the same experimental setup and neural network architectures used in the AriaNN benchmark [22]. Specifically, we evaluate FuSeFL against AriaNN using three network topologies of varying complexity:
  • Network-1: A simple multilayer perceptron composed of three fully connected layers, each followed by ReLU activation functions.
  • Network-2: A convolutional neural network (CNN) with two convolutional layers (with ReLU and max pooling) followed by two fully connected layers.
  • LeNet: A classic CNN architecture similar to Network-2 but with different layer dimensions and configurations.
All models are trained on the MNIST dataset [48], which is uniformly partitioned among all participating clients. Each client receives an equal share of the data and contributes a single vote to the federated average (i.e., FedAvg weights are uniform across clients). Training configurations and hyperparameters for each network are provided in Table IV.
TABLE IV: Training configurations for different networks
Network # Layers Dataset Learning Rate Momentum
Network1 3 MNIST 0.01 0.9
Network2 4 MNIST 0.01 0.9
LeNet 4 MNIST 0.02 0.9
Experiments are conducted on a machine equipped with an Intel(R) Xeon(R) W-2123 CPU (3.60 GHz, 4 cores). Our FuSeFL implementation is built on top of AriaNN’s implementation [22], which performs secure training using two workers while keeping the data and model private. In FuSeFL, the AriaNN training protocol is applied within each client group to collaboratively train the local model.
Figure 8: Accuracy of different federated learning schemes with four local epochs and various global epochs.
7 Evaluation Result
This section evaluate FuSeFL across multiple dimensions, including model accuracy, communication overhead, memory footprint, and training time. We use three baselines for comparison: (1) AriaNN, the original one-server, one-client secure training protocol proposed by Ryffel et al. [22], (2) AriaNN-FL, the extension of AriaNN to a federated learning setting, and (3) WW-FL, a fully offloaded secure federated learning scheme [17].
7.1 Accuracy Evaluation
Table V presents a comparison of test accuracy for FuSeFL and AriaNN when training Network1, Network2, and LeNet on the MNIST dataset, evenly partitioned across 16 clients. AriaNN achieves 98% accuracy on Network1 after 15 epochs of centralized secure training. However, extending AriaNN to a federated setting introduces accuracy degradation: with 4 global and 4 local epochs, AriaNN-FL and FuSeFL-Parallel reach 96.40%, while FuSeFL-Serial achieves 97.20%.
This accuracy gain in FuSeFL-Serial stems from its sequential intra-group training: each group first trains on one client’s data, then continues training on the second client’s data before aggregation. Aggregation is performed per group rather than per client, reducing data fragmentation.
TABLE V: Accuracy of different schemes across different networks
Scheme Epochs* Dataset Network Accuracy(%)
AriaNN[22] 15 MNIST Network1 98.00
AriaNN-FL[22] 4/4 MNIST Network1 96.40
FuSeFL-Parallel 4/4 MNIST Network1 96.40
FuSeFL-Serial 4/4 MNIST Network1 97.20
AriaNN[22] 10 MNIST Network2 98.40
AriaNN-FL[22] 3/4 MNIST Network2 97.80
FuSeFL-Parallel 3/4 MNIST Network2 97.80
FuSeFL-Serial 3/4 MNIST Network2 98.40
AriaNN[22] 10 MNIST LeNet 99.20
AriaNN-FL[22] 4/4 MNIST LeNet 99.00
FuSeFL-Parallel 4/4 MNIST LeNet 99.00
FuSeFL-Serial 4/4 MNIST LeNet 99.20
Note: (*) For the FL-based models, 4/4 means the model is trained for four global iterations and four local epochs (global/local).
Importantly, FuSeFL-Serial achieves comparable accuracy 1.43×1.43$\times$1.43 × faster than AriaNN-FL and 6.85×6.85$\times$6.85 × faster than AriaNN. For Network2 and LeNet, FuSeFL-Serial matches AriaNN’s accuracy—98.40% and 99.20%, respectively—demonstrating that FuSeFL delivers the same learning quality as traditional secure training, while offering substantial speedups due to its decentralized design. As observed in Network1, AriaNN-FL and FuSeFL-Parallel yield slightly lower accuracy for Network2 and LeNet, reaching 97.80% and 99.00%, respectively.
Accuracy Convergence Over Global Epochs: Fig. 8 shows how model accuracy evolves with increasing global training epochs for three networks (Network 1, Network 2, and LeNet). Across all models, both variations of FuSeFL (Serial and Parallel) demonstrate rapid convergence and maintain consistently high accuracy, closely tracking or outperforming AriaNN-FL. The FuSeFL-Serial variant offers a slight advantage in accuracy due to richer local training from sequential data access.
7.2 Server-Side Scalability Analysis
To assess FuSeFL’s server-side scalability, we evaluate its ability to handle an increasing number of clients under fixed hardware resources, focusing on the server’s computational overhead for training and aggregation.
As illustrated in Fig. 9, AriaNN-FL’s training time per epoch grows significantly with the number of clients, due to its centralized training coordination. For example, with 128 clients, FuSeFL-Serial achieves up to a 5.65×5.65$\times$5.65 × speedup over AriaNN-FL, while FuSeFL-Parallel achieves up to 7.83×7.83$\times$7.83 ×. The bottleneck in AriaNN-FL arises from limited parallelism: if a server has 16 compute threads, it must process clients in sequential batches, e.g., training 64 clients in four rounds of 16. This serialization leads to latency that scales linearly with client count, severely limiting scalability.
FuSeFL avoids this bottleneck by shifting training to client pairs. Servers are only responsible for secure aggregation of secret-shared model updates, not training. Since client-group training occurs in parallel and independently of server compute capacity, FuSeFL’s training time remains nearly constant even as the number of clients increases. This design enables scalable, low-latency training suitable for large-scale cross-silo deployments.
Figure 9: Training time per epoch across different schemes as the number of clients increases.
To further highlight the server-side bottleneck, Fig. 10 compares the training throughput, measured as the number of clients trained in parallel, for AriaNN-FL and FuSeFL. In AriaNN-FL, throughput initially increases with client count but quickly saturates once the number of clients exceeds the server’s available CPU cores. This plateau reflects limited server parallelism and growing serialization overhead as more clients participate. In contrast, FuSeFL decouples training from server compute constraints. Since training is offloaded to client pairs, throughput scales linearly with the number of clients, demonstrating FuSeFL’s superior scalability under high client loads.
Figure 10: Training throughput per round. FuSeFL’s throughput is not limited by the server’s computational capabilities during local training
Figure 11: Communication Overhead in FuSeFL vs. AriaNN-FL. By distributing training across client groups, FuSeFL spreads the bandwidth load across the network, substantially reducing client-to-client (C2C) communication overhead.
7.3 Communication Latency at Scale
To assess communication efficiency, we measure the total communication time for one training epoch of Network1 under AriaNN-FL and FuSeFL, assuming each client holds a dataset of 0.45 MB. We break down the overhead into two channels: client-to-client (C2C) and server-to-client/client-to-server (S2C/C2S), as shown in Fig. 11.
In AriaNN-FL (Fig. 11(a)), the server is actively involved in training each client’s model via secure computation, requiring bidirectional communication with all clients per epoch. As the number of clients increases, this centralized setup leads to a linear increase in both S2C/C2S and C2C communication. With 100,000 clients, the overall communication overhead exceeds 11,000 seconds, driven by server-side congestion and sequential communication demands. Even under high parallelism, the server remains a bottleneck due to the required per-client secure coordination.
FuSeFL, shown in Fig. 11(b), mitigates this issue by fully decentralizing the training to client pairs. In this design, C2C communication is localized and parallelized across the network. As a result, this overhead remains nearly constant regardless of scale (e.g.,  0.21 seconds even at 10,000 clients), as communication occurs simultaneously within independent groups. S2C/C2S communication in FuSeFL, which only involves the transmission of final model shares to the servers for secure aggregation, scales linearly but at half the rate of AriaNN-FL. This is because only one model per group is uploaded per round, cutting the volume of communication in half. For instance, at 10,000 clients, FuSeFL’s total communication time is only 45.21 seconds compared to 1,123 seconds in AriaNN-FL, yielding a 25× reduction.
This efficiency gain highlights FuSeFL’s ability to scale securely and efficiently across thousands of clients in cross-silo settings, where communication bandwidth and latency are critical deployment constraints.
7.4 Memory Overhead
Aggregator nodes in federated learning must store multiple instances of local model parameters until all client updates are received for weighted aggregation at the end of each epoch [36]. This requirement leads to considerable memory overhead, especially as the number of clients increases. Figure 12 shows the server-side memory usage for storing model instances. In conventional FedAvg, the server cannot incrementally aggregate updates and must retain all client updates simultaneously, causing memory consumption to scale linearly with the number of participants.
FuSeFL improves this by organizing training into client pairs, each producing a single aggregated model update per round. As a result, the number of model instances stored on the server is effectively halved. For instance, with 128 clients, AriaNN-FL stores 128 local models plus the global model. FuSeFL reduces this to 64 local models while preserving model quality and reducing memory load. This improvement is especially beneficial in cross-silo deployments, where the number of participating institutions can be large and model sizes (e.g., VGG-16 at 528MB in FP32) significantly increase server storage demands.
Figure 12: Memory footprint of model instances at the server. FuSeFL reduces server-side memory usage by 50% compared to AriaNN-FL by offloading model training to client groups, thereby minimizing the number of model instances the server must store.
7.5 Comparison with WW-FL[17]
To compare FuSeFL with WW-FL [17], we configure both systems to use AriaNN as their underlying MPC protocol and evaluate them under identical training parameters shown in Table VI. Notably, FuSeFL performs local training without relying on third-party computation servers, whereas WW-FL requires external servers within each cluster to train secret-shared client data.
TABLE VI: Training parameters of WW-FL and FuSeFL
Parameter WW-FL FuSeFL
# clients 100 100
# clients per round 100 100
# MPC clusters 10 N/A
# clients per clusters 10 N/A
# clients per clusters per round 10 N/A
Size of clients dataset (MB) 1 1
# Servers used for aggregation 2 2
# Servers used for training 20 0
Note: N/A means FuSeFL does not have the concept of clusters.
Table VII compares the communication cost per training round. In WW-FL, each client must offload both secret shares of its data to designated servers for secure computation, doubling the communication overhead. In FuSeFL, each client retains one share locally and sends the other to its group partner, reducing total communication by 50%. Additionally, the critical communication path, defined as the most bandwidth, congested channel—is substantially lighter in FuSeFL (2 MB vs. 10 MB), improving scalability.
TABLE VII: Comparison of FuSeFL with WW-FL for one round of training
Parameter WW-FL FuSeFL
Total uploaded data (MB) 200 100
Uploaded data per client (MB) 2 1
Critical communication path (MB) 10 2
WW-FL introduces a trade-off between model accuracy and communication overhead. As the number of MPC clusters decreases, each cluster includes more clients, increasing model accuracy but requiring more data to be transferred per client. Table VIII shows that accuracy improves from 93.00% (C=16) to 96.30% (C=4), but data transfer rises from $N$ to $4N$ per client, undermining FL’s goal of minimizing communication with centralized servers.
In contrast, FuSeFL achieves 95.10% accuracy with only $2N$ data transfer, matching WW-FL with $C=8$ but with no third-party servers required. Moreover, as illustrated in Fig. 8, FuSeFL can achieve further accuracy gains by increasing global epochs without increasing the communication burden.
TABLE VIII: Comparison of FuSeFL with WW-FL in terms of accuracy and communication
WW-FL FuSeFL
C=16 C=8 C=4
Accuracy 93.00 95.10 96.30 95.10
User Data Transfer N 2N 4N 2N
Unlike WW-FL, which assumes trusted computation servers and requires offloading full data shares for training, FuSeFL retains one share locally and sends only the other to a peer, eliminating the need for a trusted zone. This not only strengthens data privacy but also simplifies deployment in real-world cross-silo settings. Furthermore, by decentralizing secure training to the client layer, FuSeFL avoids WW-FL’s server-heavy design—drastically reducing infrastructure costs, minimizing inter-zone latency, and enabling more scalable, lightweight deployments.
7.6 Robustness Under Threat Model
This section analyzes the robustness of FuSeFL under the adversarial threat model described in Section 3. The primary security goals are to protect the confidentiality of user data, model parameters, and intermediate updates in the presence of semi-honest adversaries. We consider two prominent threat vectors: unauthorized observation and model theft.
Protection Against Unauthorized Observation. Adversaries may attempt to extract sensitive information by intercepting communication between clients or by observing model parameters during training. In FuSeFL, all private values, including client data, model parameters, and gradients, are secret-shared using additive masking. These shares appear statistically indistinguishable from random noise, preventing inference even in the case of eavesdropping.
To securely perform training, FuSeFL employs the AriaNN protocol [22], which guarantees confidentiality through function sharing and additive masking. During secure training, no single party observes plaintext data or model parameters. Additionally, each client retains one share of its data locally and never reveals its complete input to its training partner, preserving local privacy. These guarantees are backed by established cryptographic protocols validated in prior literature [19, 43, 45, 39, 40].
Protection Against Model Theft. Model confidentiality is a critical requirement in federated settings, particularly when the model itself represents proprietary intellectual property. In FuSeFL, the global model is secret-shared between two non-colluding servers and then transmitted to clients as encrypted shares. Each client receives only one share and cannot reconstruct the full model independently. Even if an adversary intercepts a share during transmission, it reveals no meaningful information.
Furthermore, aggregation servers do not have access to the full model updates; they operate only on shares during the FedAvg procedure. These protections collectively ensure that no party, client or server, can exfiltrate the full model under the semi-honest assumption. FuSeFL thereby mitigates the risk of model theft during both training and communication phases.
8 Related Work
To protect privacy in FL, several cryptographic and statistical methods have been proposed. Homomorphic encryption (HE) allows computation over encrypted data but is computationally intensive, especially in cross-silo settings with large-scale datasets [3, 4]. Differential privacy (DP) introduces calibrated noise to gradients or updates to provide formal privacy guarantees [5, 6], but this often comes at the cost of reduced model accuracy. Secure multi-party computation (MPC), in contrast, offers strong privacy guarantees without degrading accuracy, though it typically incurs high communication costs.
MPC-based approaches have gained traction in secure FL. SecureML [19], SecureNN [20], Falcon [21], and AriaNN [22] propose secure training protocols using various cryptographic primitives, including additive secret sharing [40, 39], function secret sharing [43, 45], and Yao’s circuits [49]. These works typically offload both data and models to trusted third-party servers for secure computation. However, they assume trust in infrastructure and require clients to fully transmit their data shares, increasing bandwidth demands and collusion risks.
WW-FL [17] extends this idea by introducing a cluster-based MPC training framework. Clients are grouped into clusters and offload secret-shared data to server pools, which perform collaborative training. While this improves server-side load distribution, it still assumes trusted execution on third-party servers and leads to high communication overhead as the number of clients or cluster sizes increase. Additionally, WW-FL relies on a client-side trusted zone for safely outsourcing data, which may not be feasible in practical deployments.
Other complementary works attempt to establish model ownership through watermarking [15, 16]. Backdoor watermarking uses a trigger dataset, while parameter watermarking embeds ownership directly in model weights. While effective for post-training verification, these methods do not provide confidentiality during training and are orthogonal to secure computation.
Finally, hierarchical FL [50, 51, 17] reduces communication costs by introducing intermediate aggregators for groups of clients. This approach improves scalability and bandwidth usage but does not inherently address the confidentiality of models or data during training.
9 Conclusion
In this work, we introduced FuSeFL, a fully secure and scalable federated learning framework tailored for cross-silo environments. By decentralizing training across client pairs using lightweight MPC and assigning a minimal aggregation role to the servers, FuSeFL eliminates central bottlenecks and ensures complete confidentiality of client data, model parameters, and training updates. Our design not only mitigates key inference threats, including model inversion and gradient leakage, but also significantly reduces communication latency and server resource demands. Empirical evaluations show that FuSeFL achieves strong accuracy and performance improvements over existing secure FL solutions, establishing it as a practical and robust alternative for secure collaborative learning in regulated domains.